OpenStudy (anonymous):

Please Help!! Which set of integers is a Pythagorean triple? A. 2, 3, 6 B. 3, 4, 5 C. 4, 5, 9 D. 5, 7, 12

OpenStudy (anonymous):

pretty sure its a

OpenStudy (anonymous):

pythagoras \[a^2 + b^2 = c^2\] biggest number is always c^2. check, if putting the other two numbers satisfy the equation.

OpenStudy (anonymous):

So what is the answer?

OpenStudy (solomonzelman):

\[a)~~~~2^2+3^2=6^2\] \[b)~~~~3^2+4^2=5^2\] \[c)~~~~4^2+5^2=9^2\] \[d)~~~~5^2+7^2=12^2\] one of the statements is true and therefore a Pythagorean triple. Which one?

OpenStudy (anonymous):

its 3,4,5